Compute the following ratios. Do not round your calculations. (Always use cell references and formulas where appropriate to receive full credit.) 2019 2018 Working Capital Current Ratio Acid-Test (Quick) Ratio Cash Ratio Accounts Receivable Turnover Days' Sales in Receivables Inventory Turnover Days' Sales in Inventory Gross Profit Percentage Debt Ratio Debt to Equity Ratio Times-Interest- Earned Ratio Profit Margin Ratio Rate of Return on Total Assets Asset Turnover Ratio Rate of Return on Common Stockholders' Equity Earnings per Share (EPS) Price/Earnings Ratio Dividend Yield Dividend Payout HINTS
HINTS Cell Hint: C5 | Begin the formula with an C7 Use the function SUM) to calculate the numerator. c9 (0.75*Vertical Analysis'!C8)/(('Horizontal Analysis'!c14+Horizontal Analysis'!D14)/2) C10 | Assume 365 days in a year and use the correct ratio in this worksheet for your calculations. use the balance in the horizontal analysis worksheet for your calculations, and press the enter key. equals sign (),
June 30 2018 and 2019 5 6 Increase (Decrease) Amount 7 Percentage 2019 2018 8 Assets 9 10 Current Assets $ 125,000,00 $ 685,000.00 $ $ 5.04% 5.38% Cash 119.000.00 6.000.00 11 $ 650.000.00 Short Term Investments 35,000.00 12 27,000.00 $ (5,000.00) Accounts Receivable 225,000.00 198.000.00 13.64% 13 Merchandise Inventory S S -7.14% 65.000.00 70,000.00 14 4,000.00 S 67.000.00 S 43.000.00 Other Current Assets 195,000.00 191,000.00 S 2.09% 15 16 Total Current Assets 17 PPE 18 Total Assets $1,295,000.00 $ 875,000.00 $2,170,000.00 $2,060,000.00 $1,228,000.00 $ 832.000,00 5.46% 5.17% $110,000.00 5.34% 19 Liabilities 20 21 Current Liabiltiies: Accounts payable Accrued Liabilities 265,000.00 251,750.00 641,000.00 $ 725.523.00 906,000.00 977,273.00 S 13,250.00 (84,523.00) S (71,273.00) 5.26% 22 -11.65% 23 24 Total Current Liabilities -7.29% 25 Long-Term Liabilities Bonds Payable Mortgage payable 28 Total Long Term Liabilities 29 Total Liabilities $ 100,000.00 $ (25,000.00) $ 75,000.00 3,727.00 250,000.00 150,000.00 $150,000.00 175,000.00 400,000.00 325,000.00 $1,306,000.00 $1,302,273.00 66.67% 26 -14.29% 27 23.08% S 0.29% Stockholders' Equity 30 225,000.00 $ 225,000.00 31 Common Stock S 0.00% 32 Paid in Capital 33 Reatined Earnings 34 Total Stockholders' Equity 35 Total Liabilities & Stockholders' Equity S 58,000.00 581,000.00 58.000.00 S 0.00% $ 474,727.00 22.39% $106,273.00 $106,273.00 $110,000.00 S 864,000.00 $2,170,000.00 $2,060,000.00 $ 757,727.00 14.03% 5.34% 36
Riverside Sweets Income statement Year ended June 30,2019 A P $28,00,000 Net Sales 100% Cost of goods sold 1551600 55% Gross Profit 1248400 45% Operating expenses 450540 16% 797860 28% Operating Income Other Income and (Expenses) Interest expenses 15000 1% 782860 Income Before Income taxes 28% Income Tax expenses 153529 5% $6,29,331 Net Income 22%
Answer #1

1.Working Capital = Current Assets - Current Liabilities

2019 = 1295000 - 906000 = 389000 ; 2018 = 1228000 - 977273 = 250727

2.Current Assets = Current Assets / Current Liabilities

2019 = 1295000 / 906000 = 1.43 ; 2018 = 1228000 / 977273 = 1.26

3.Acid Test Ratio = Quick Assets / Current Liabilities

2019 = ( 125000 + 685000 + 225000 ) / 906000 = 1.14 ; 2018 = ( 119000 + 650000 + 198000 ) / 977273 = 0.989

4.Cash Ratio = Cash / Current Liabilities

2019 = 125000 / 906000 = 0.138 ; 2018 = 119000 / 977273 = 122

5.Accounts Receivable Turnover = Sales / Avg.Receivables

2019 = 2800000 / ( 225000 + 198000 ) / 2 = 2800000 / 211500 = 13.24

6.Days Sales in Receivables = 365 / Accounts Receivable Turnover = 365 / 13.24 = 27.57 Days

7.Inventory Turnover = Cost of goods sold / Avg.Inventory = 1551600 / ( 65000 + 70000 ) / 2 = 1551600 / 67500 = 22.99

8.Days Sales in Inventory = 365 / Inventory Turnover = 365 / 22.99 = 15.88 Days

9.Gross Profit Percentage = Gross Profit / Sales = 1248400 / 2800000 = 44.59%

10.Debt Ratio = Total Liabilities / Total Assets

2019 = 1306000 / 2170000 = 0.60 ; 2018 = 1302273 / 2060000 = 0.63

11.Debt to Equity Ratio = Total Liabilities / Total Equity

2019 = 1306000 / 864000 = 1.51 ; 2018 = 1302273 / 757727 = 1.72

12.Times Interest Earned Ratio = Operating Income / Interest Expense

2019 = 797860 / 15000 = 53.19

13.Profit Margin Ratio = Net Income / Sales = 629331 / 2800000 = 22.48%

14.Rate of Return on Total Assets = Net Income /Total Assets = 629331 / 2170000 = 29%

15.Asset Turnover Ratio = Sales / Avg. Total Assets = 2800000 / ( 2170000 + 2060000 ) / 2 = 2800000 / 2115000 = 1.32

16.Rate of Return on Common Stockholders Equity = Net Income / Avg. Common Stockholders Equity

629331 / ( 864000 + 757727 ) / 2 = 629331 / 810864 = 0.776

17.Earning per share = Net Income / Outstanding no. of shares = 629331 / 225000/10 = 27.97

18.Price Earnings Ratio = Market Price / EPS ...Market price not provided

Dividend not provided